You have been hired by Kia as manager for its Pakistan operations. Assume following is the short￾run production function at their assembly plant outside Karachi:

Q = 10L2 – 0.5 L3

 where L is variable input labor, Q is output of Cars assembled


b. Demonstrate the relationship between Total Production, Marginal Product and 

AverageProduct in a hypothetical graph and clearly label the three stages as per the 

values of Lyou observed in 


1
Expert's answer
2021-01-29T11:16:02-0500

Part A

"Q" "=" "10L""^2" "-0.5L^3"

Average Product, "AP=" "Q\/L"

"10L^2-0.5L^3\/L"

"10L-0.5L^2"

At maximization we equate to zero & differentiate with respect to L.

"d(AP)\/dL=0"

"10-1.5*2L=0"

"L=10"

Stage 1 is as a result of average product increase. The range is from L=0 to L=10

Marginal product,

"MP=dQ\/dL"

"MP=10*2L-0.5*3L^2"

"MP=20L-1.5L^2"

At Stage 2 margin product decreases but still remains positive.

At Stage 3 marginal product is equal to zero

"MP=0"

"20L-1.5L^2=0"

"20=1.5L"

"L=20\/1.5"

"L=13.33"

Part B